OpenStudy (anonymous):

Find the derivative of the function. y=sqrt( 3x+sqrt( 3x+sqrt( 3x)))

OpenStudy (dumbcow):

nested radical function... you have to use chain rule let me try to simplify it using substituted functions \[u = \sqrt{3x} \rightarrow du = \frac{3}{2 \sqrt{3x}}\] \[v = \sqrt{3x + \sqrt{3x}} = \sqrt{3x +u}\] \[\rightarrow dv = \frac{3 + du}{2 \sqrt{3x+u}}\] \[y = \sqrt{3x +v} \rightarrow dy = \frac{3 + dv}{2 \sqrt{3x +v}}\] now you have dy, just substitute everything in